Could you help me to solve the equation:
16/x^2=8/x+6
I know the LCD is x^2(x+6) and I multiply each term and I get
16(x+6)=8x^2
which turns into the equation
8x^2-16x+96=0 (right?)
which is where I get confused. I'm not sure if I am suposed to factor it out or use the quadratic formula to solve it. I know though that the answer involves a square root.
Please help!

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

16/x^2=8/x+6
I know the LCD is x^2(x+6) and I multiply each term and I get
16(x+6)=8x^2
which turns into the equation
8x^2-16x+96=0 (right?)
which is where I get confused.
--------------------------------
16/x^2=8/x+6
I know the LCD is x^2(x+6) and I multiply each term and I get
16(x+6)=8x^2
which turns into the equation
8x^2-16x-96=0
x^2 - 2x - 12 = 0
x = [2 +- sqrt(4-4*-12)]/2
x = [2 +- sqrt(52)]/2
x = [2 +- 2sqrt(13)]/2
x = 1 + sqrt(13) or x = 1-sqrt(13)
=========================
Cheers,
Stan H.

Answer by scott8148(6628)

the 96 should be negative __ dividing by 8 __ x^2-2x-12=0

not factorable, so use quadratic formula __ x=[2 +- sqrt(4+48)]/2 __ x=1 +- sqrt(13)